# ZeroDL: Zero-shot Distribution Learning for Text Clustering via Large Language Models

### v1

@@ Abstract
The recent advancements in large language models (LLMs) have brought significant progress in solving NLP tasks. Notably, in-context learning (ICL) is the key enabling mechanism for LLMs to understand specific tasks and grasping nuances. In this paper, we propose a simple yet effective method to contextualize a task toward a specific LLM, by (1) observing how a given LLM describes (all or a part of) target datasets, i.e., open-ended zero-shot inference, and (2) aggregating the open-ended inference results by the LLM, and (3) finally incorporate the aggregated meta-information for the actual task. We show the effectiveness of this approach in text clustering tasks, and also highlight the importance of the contextualization through examples of the above procedure.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performances on various downstream tasks Devlin et al. (2019); Radford et al. (2019). These also exhibit the ability to understand the context of input text, known as in-context learning (ICL) Brown et al. (2020); OpenAI (2023). ICL allows to leverage LLMs for specific tasks without further extensive training. However, effective use of ICL hinges on well-designed prompts.
While prompts with few-shot examples can demonstrably improve performance, they can easily overfit a model to the examples (Perez et al. (2021); Mizrahi et al. (2023); inter alia). This led to a growing interest in zero-shot learning, which reduces the need for intricate few-shot selection. Recent advancements in zero-shot learning involve incorporating more sophisticated use of prompt structures, such as Chain-of-Thought Wei et al. (2022), zero-shot reasoning Kojima et al. (2022), and models trained to follow instructions Ouyang et al. (2022); Chung et al. (2024).
Figure 1: Illustration of proposed method ZeroDL. While in-context learning (ICL) relies on examples tailored to specific tasks, ZeroDL aggregates all the outputs from these zero-shot inferences, resulting in meta-level information. This meta-level information is then used by the LLM to generate its final predictions.
Motivated by the core principle of ICL–providing task and data contexts within prompts–we propose an approach to construct more effective zero-shot prompts by understanding how LLMs describe datasets across prompting outputs. As illustrated in Figure 1, the method, called Zero-shot Distribution Learning (ZeroDL), first aims to learn data distributions through zero-shot inferences. The method comprises two key components: (1) open-ended zero-shot inference, and (2) output aggregation. A zero-shot prompt is then constructed with the generated meta information, and used for actual task. This method takes advantage of the self-generated frame of LLMs to successfully carry out a given task.
We exemplify the effectiveness of ZeroDL with text clustering tasks, where we assume an absence of ground-truth class labels. Unlike traditional clustering algorithms such as K-Means representing groups based on distance metrics in embedding space, our method works in a text-to-text format. Hence, our approach allows clustering with perspectives and/or specific context. For instance, the sentences "I love this movie" and "I hate this movie" express opposite sentiment but belong to the same topic cluster of movie reviews. Refer to Table 8 in the Appendix for potential risk of absence of the perspectives.
@@
  • ZeroDL enables models to perform text-based clustering, empowering them to handle data with specific context, which offers advantages over embedding-based clustering methods.
  • Our experiments show that ZeroDL is competitive against embedding-based clustering methods on several datasets. ZeroDL even achieves better performances than models with ground-truth class labels in some cases.
@@ 2 Related Works
While there have been a few attempts to leverage LLMs for clustering tasks, majority of existing approaches rely on traditional methods like K-Means clustering based on LLM-generated embeddings Petukhova et al. (2024); BehnamGhader et al. (2024). In contrast, our approach leverages text-level prompting, which injects specific viewpoints into LLMs, enabling it to perform more targeted and contextualized clustering.
@@
The importance of appropriate ground-truth labels extends beyond clustering and permeates ICL. While Min et al. (2022) observed cases where the input-label correspondence does not play significant roles, Yoo et al. (2022) argued that the impact heavily depends on target tasks and experiment settings. Compared with the few-shot settings these works assumed, our method is with zero-shot for clustering. We believe that this work would serve as a reference that supports the appropriate class information is crucial for the performances, as we develop ZeroDL to leverage the power of generative LLMs to propose effective class labels, i.e., adjusting a task from LLMs’ perspective.
@@
TF-IDF KMeans 52.1 53.0 26.2 35.2 43.8 55.9 62.8 44.7 46.5 46.7 48.8
mistral Gold(C-T) 87.5 77.0 41.8 50.8 60.7 74.2 85.2 40.8 62.5 64.5 68.1
-7b-inst Gold(T-C) 91.7 82.5 43.3 51.8 82.7 84.1 82.7 50.9 73.8 71.5 72.7
ZeroDL(C-T) 85.3 82.4 43.4 51.6 75.4 64.0 73.3 47.2 69.4 65.8 65.1
ZeroDL(T-C) 92.2 79.4 36.0 48.2 81.0 62.2 78.4 51.8 73.4 67.0 66.4
llm2vec+KMeans 62.1 55.2 30.3 56.0 84.4 96.1 70.8 48.0 46.2 61.0 67.5
Llama-3 Gold(C-T) 66.5 73.7 42.1 48.9 55.9 71.3 75.5 35.2 54.7 58.2 61.1
-8b-inst Gold(T-C) 93.2 84.2 46.0 53.5 70.2 74.6 79.2 49.3 75.4 69.5 70.3
ZeroDL(C-T) 89.8 74.2 42.6 37.6 69.0 65.9 52.5 46.8 48.2 58.5 56.4
ZeroDL(T-C) 66.7 79.1 41.1 49.2 80.4 73.2 63.7 49.1 48.6 61.2 60.8
llm2vec+KMeans 53.2 53.2 29.6 54.8 85.1 94.5 72.7 47.4 50.4 60.1 66.2
Llama-3 Gold(C-T) 94.4 88.9 53.0 56.6 83.2 94.8 95.8 65.3 82.4 79.4 80.7
-70b-inst Gold(T-C) 95.2 89.4 52.3 58.2 81.1 95.8 91.0 63.8 82.6 78.8 80.3
ZeroDL(C-T) 94.3 90.0 42.7 47.7 83.3 55.3 35.4 51.2 82.0 64.7 57.5
ZeroDL(T-C) 95.0 89.4 41.5 49.3 72.6 47.0 36.1 53.2 81.4 62.8 56.2
@@
Table 1: The performance of LLMs for text clustering. C-T denotes the prompt order with class information then input text. T-C is the reversed. Bold means the best accuracy and underline the outperforming cases than ground-truth (i.e., Gold) class label setting. Note that llm2vec for Llama3-70b is not publicly available at the moment.
3 Proposed Method: ZeroDL
The framework is illustrated in Figure 1.
[…]
Stage 2: Aggregation. The open-ended predictions lack constraints, leading to potentially inconsistent output formats. For instance, the model might predict "positive" and "non-positive" classes, while the ground-truth is "positive" and "negative". The predictions could even be entire sentences. To address the inconsistencies in the open-ended predictions, we employ aggregation strategy.
Before the aggregation, we count the frequency of each predictions and sort it. After that, the predictions which frequency is only 1 are dropped in order to filter out extraordinary predictions and save computation. Next, we iteratively construct subsets of the predictions by removing the least frequent predictions one by one. This process results in a list of subsets, denoted as [ $\{\hat{m_{1}},\hat{m_{2}},\cdots,\hat{m_{U}}\}$, $\{\hat{m_{1}},\hat{m_{2}},\cdots,\hat{m_{U-1}}\}$, $\cdots$ $\{\hat{m_{1}}\}$ ] where $U$ denotes the number of remained predictions. We input all the subset to provide more weights to the frequently occurred predictions. The prompt for aggregation is as follows:
[…]
While the prompt instructs the LLM to predict a specific number of classes ([NUM_CLUSTER_CLASS]), the model outputs often still lack coherence, especially in generating exact number of classes. To address this, we select the LLM outputs where the number of aggregated classes matches the pre-defined number of cluster classes. We then use the most frequent class pairs aggregated from prompt (sub)sets as meta-information $\hat{M}$. This information represents the model’s understanding of potential views over the entire data.
Stage 3: Leveraging Meta-Information for Prediction. We incorporate the aggregated meta-information ($\hat{M}$) into the classification prompt [^1] The order of Text: and Class description: can be reversed. to enhance the model’s prediction capabilities:
[…]
- Class 0: $\hat{M_{0}}$
- Class 1: $\hat{M_{1}}$
- Class $\cdots$: $\hat{M_{\cdots}}$
- Class $C$: $\hat{M_{C}}$
[…]
the text to the best [type_of_task] class.
Within the prompt, we include placeholders denoted as - Class n:. These placeholders serve as templates for the LLM to parse.
By incorporating the meta-information ($\hat{M}$) into the prompt, we enable the LLM to perform conditioned classification within the clustering context. The generated meta-information might not directly correspond to the ground-truth classes, potentially leading to variations in performance.
4 Experiments
Models. Our experiments are conducted on mistral-7b-instruct-v0.2 Jiang et al. (2023) due to its impressive performance even with a relatively small parameter size compared with other LLMs. We additionally report results obtained with the Llama-3 family models Meta (2024) from HuggingFace Wolf et al. (2020) ran with vLLM library Kwon et al. (2023). Greedy decoding is employed for generation [^2] Note that the current version of vLLM might exhibit non-deterministic behavior. and the final prediction is done with guided_choice hyperparameters. We follow the official instruction formatting guidelines. The results are averaged from 5 runs. We also tested llm2vec BehnamGhader et al. (2024) with KMeans approach where applicable.
Datasets and Evaluation. We use 6 datasets: IMDB Maas et al. (2011), SST-2, SST-5 Socher et al. (2013), YelpReivews Zhang et al. (2015) for sentiment classification and AGNews, DBpedia Lehmann et al. (2015), YahooAnswers Chang et al. (2008) for topic classification. Further information is presented in Table 4 of the Appendix.
@@ 4 Experiments
Table 2: The example of generated class labels in 5-class sentiment classification (SST-5), and topic classification (AGNews). ZeroDL can generates alternative class labels and its description. Additional examples are in Table 7.

### v2

@@ Abstract
The advancements in large language models (LLMs) have brought significant progress in NLP tasks. However, if a task cannot be fully described in prompts, the models could fail to carry out the task. In this paper, we propose a simple yet effective method to contextualize a task toward a LLM. The method utilizes (1) open-ended zero-shot inference from the entire dataset, (2) aggregate the inference results, and (3) finally incorporate the aggregated meta-information for the actual task. We show the effectiveness in text clustering tasks, empowering LLMs to perform text-to-text-based clustering and leading to improvements on several datasets. Furthermore, we explore the generated class labels for clustering, showing how the LLM understands the task through data.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performances on various downstream tasks Devlin et al. (2019); Radford et al. (2019). These also exhibit the ability to understand the context of input text, known as in-context learning (ICL) Brown et al. (2020); OpenAI (2023). ICL allows leveraging LLMs for specific tasks without further extensive training. However, effective use of ICL hinges on well-designed prompts.
While prompts with few-shot examples demonstrably improve performance, they can easily overfit a model to the examples (Perez et al. (2021); Mizrahi et al. (2023); inter alia). This led to a growing interest in zero-shot learning, which reduces the need for intricate few-shot selection. Recent advancements in zero-shot learning involve incorporating more sophisticated use of prompt structures, such as Chain-of-Thought Wei et al. (2022), zero-shot reasoning Kojima et al. (2022), and models trained to follow instructions Ouyang et al. (2022); Chung et al. (2024). However, how to design prompts for target tasks remains challenging.
Figure 1: Illustration of proposed method ZeroDL. While in-context learning (ICL) relies on examples (D) tailored to specific tasks, ZeroDL aggregates all the outputs from these zero-shot inferences ($\hat{m}$), resulting in meta-level information ($\hat{M}$). This information is then used by the LLM to generate its final predictions.
Motivated by the core principle of ICL–providing task and data contexts within prompts–we propose an approach to construct more effective zero-shot prompts by understanding how LLMs describe datasets across prompting outputs.
As illustrated in Figure 1, Zero-shot Distribution Learning (ZeroDL) aims to learn data distributions through zero-shot inferences. The method comprises two key components: open-ended zero-shot inference and output aggregation. Zero-shot prompts are then constructed with the generated meta information, and used for actual task. This method takes advantage of the self-generated frame of LLMs to successfully carry out a given task.
We exemplify the effectiveness of ZeroDL on text clustering tasks where complete task descriptions cannot be provided due to an absence of ground-truth class labels. In addition, our method works in a text-to-text format, allowing clustering with specific context. For instance, "I love this movie" and "I hate this movie" express opposite sentiment but belong to the same cluster of movie reviews (see Table 9 in the Appendix for potential risk of absence of the perspectives).
@@
  • ZeroDL allows models to perform text-based clustering, empowering them to handle data with specific context, which offers advantages over embedding-based clustering methods.
  • ZeroDL is competitive against embedding-based clustering methods on several datasets. Notably, ZeroDL even achieves better performances than models with ground-truth class labels in some cases.
@@ 2 Related Works
The majority of existing LLM-based clustering approaches rely on traditional methods like K-Means, employing LLM-generated embeddings as input Petukhova et al. (2024); BehnamGhader et al. (2024). ClusterLLM Zhang et al. (2023) introduced LLM-guided refinement of embedding-based clustering models. In contrast, our approach ZeroDL, directly leverages text-level prompting to inject specific viewpoints into LLMs, enabling more targeted and contextualized clustering.
Other works have explored using LLMs as clustering models Wang et al. (2023); Viswanathan et al. (2023); Pham et al. (2023); Huang and He (2024), but these approaches typically relied on few-shot settings with access to gold labels. IDAS De Raedt et al. (2023) addressed this limitation by selecting representative data using an auxiliary embedding model to generate pseudo-labels without demonstrations. These labels are subsequently refined through in-context learning.
ZeroDL distinguishes itself by operating in a fully zero-shot setting, eliminating the need for additional embedding models. Moreover, while IDAS’s class labels are primarily influenced by a limited set of data points and generate the labels independently, ZeroDL considers the entire data distribution through an aggregation step. This enables our method to generate class labels in an auto-regressive manner, taking into account the relationships between different labels.
@@
The importance of appropriate ground-truth labels extends beyond clustering and permeates ICL. While Min et al. (2022) observed cases where the input-label correspondence does not play significant roles, Yoo et al. (2022) argued that the impact heavily depends on target tasks and experiment settings. We believe that this work would serve as a reference that appropriate class labels can be successfully generated by LLMs themselves.
3 Proposed Method: ZeroDL
Stage 1: Open-Ended Inference
We begin by designing a prompt for zero-shot classification. This prompt intentionally avoids any detailed information about the task, minimizing the risk of overfitting. Based on the idea, we opt for the simplest prompt format:
Open-Ended Inference Prompt Template Text: [text] Classify the text to the best [type_of_task] class.
where [text] is the input data and [type_of_task] provides view of the task. In the experiment, it can be either sentiment or topic. Leveraging this prompt, we perform model inferences on all the input data ($D$). This process generates open-ended class predictions, which will be denoted as $\hat{m}$.
Stage 2: Aggregation
The open-ended predictions lack constraints, leading to potentially inconsistent output formats. For instance, the model might predict "positive" and "non-positive" classes, while the ground-truth is "positive" and "negative". The predictions could even be entire sentences. To address the inconsistencies in the open-ended predictions, we employ aggregation strategy.
Before the aggregation, we count the frequency of each predictions and sort it, denoted as $\hat{m}_{1},\hat{m}_{2},\cdots,\hat{m}_{U}$ where $U$ denotes the unique number of predictions and the frequency of $\hat{m}_{n}$ is equal or larger than $\hat{m}_{n+1}$. After that, the predictions which frequency is only 1 are dropped in order to remove extraordinary predictions and save computation. Next, we iteratively construct subsets of the predictions by removing the least frequent predictions one by one. This process results in a list of subsets, denoted as $S$; $S_{U}=\{\hat{m}_{1},\hat{m}_{2},\cdots,\hat{m}_{U}\}$, $S_{U-1}=\{\hat{m}_{1},\hat{m}_{2},\cdots,\hat{m}_{U-1}\}$, $\cdots$, $S_{1}=\{\hat{m}_{1}\}$. We input each subset to the LLM, providing more weights to the frequently occurred predictions (e.g., $\hat{m}_{1}$ is repeatedly included in the subsets while $\hat{m}_{U}$ occurs only once). The prompt template for aggregation is as follows:
Aggregation Prompt Template [type_of_task] List: $S_{U}$ $\hat{m_{1}}$ $\hat{m_{2}}$cdots $\hat{m}_{U-1}$ $\hat{m_{U}}$ $S_{U-1}$ $\hat{m}_{1}$ $\hat{m}_{2}$cdots $\hat{m}_{U-1}$ $S_{\cdots}$ $\hat{m}_{1}$ $\hat{m}_{2}$cdots $S_{1}$ $\hat{m}_{1}$ Aggregate the [type_of_task] List into [NUM_CLUSTER_CLASS] classes.
where [NUM_CLUSTER_CLASS] is pre-defined number of classes to cluster. However, the model outputs often still lack coherence, especially in generating exact number of classes. To address this, we select the generation results only when the number of aggregated classes matches the pre-defined number of cluster classes. We then use the most frequent aggregated classes as meta-information $\hat{M}$ (class labels in the experiments). This information represents the model’s understanding of potential views over the entire data.
Stage 3: Leveraging Meta-Information
We incorporate the aggregated meta-information ($\hat{M}$) into the original prompt [^1] The order of Text: and Class description: can be reversed.:
Final Prediction Prompt Template Text: [text]
Class description:
- Class 0: $\hat{M_{0}}$ Class $\cdots$: $\hat{M_{\cdots}}$ Class $C$: $\hat{M_{C}}$ Based on the class description, classify the text to the best [type_of_task] class.
@@ 3 Proposed Method: ZeroDL
hcx-seed ZeroDL(C-T) 56.2 62.5 33.6 43.6 33.1 24.1 16.9 21.3 18.3 34.4 33.1
-0.5b-it IDAS(C-T) 88.3 74.9 27.7 35.0 29.4 17.0 26.8 - - - -
Gold(C-T) 72.1 75.2 35.2 27.6 31.4 31.2 24.1 18.3 18.3 37.0 33.6
ZeroDL(T-C) 86.7 72.2 36.0 43.9 59.5 32.3 44.5 26.7 22.8 47.2 46.4
IDAS(T-C) 89.4 62.9 28.9 28.7 27.3 23.7 27.2 - - - -
Gold(T-C) 89.1 73.7 34.4 39.1 59.4 61.1 57.2 40.3 56.2 56.7 56.3
mistral llm2vec+KMeans 62.1 55.2 30.3 56.0 84.4 96.1 70.8 48.0 46.2 61.0 67.5
-7b-it ZeroDL(C-T) 88.8 85.7 40.5 48.7 75.2 58.9 64.1 47.2 61.4 63.4 61.5
IDAS(C-T) 51.1 63.1 41.4 37.3 38.1 32.1 21.8 30.2 41.8 39.7 35.4
Gold(C-T) 87.5 77.0 41.8 50.8 60.7 74.2 85.2 40.8 62.5 64.5 68.1
ZeroDL(T-C) 90.2 84.2 36.0 46.8 79.5 56.7 72.2 51.0 66.3 64.8 63.0
IDAS(T-C) 50.1 68.1 39.6 41.0 44.6 35.8 27.6 33.3 46.1 42.9 38.9
Gold(T-C) 91.7 82.5 43.3 51.8 82.7 84.1 82.7 50.9 73.8 71.5 72.7
Qwen-2.5 ZeroDL(C-T) 87.6 74.8 45.5 53.8 70.9 62.3 75.9 45.3 68.8 65.0 65.8
-7b-it IDAS(C-T) 93.3 87.3 42.7 51.1 68.8 47.4 68.4 30.3 41.8 58.9 59.0
Gold(C-T) 84.3 87.4 45.0 49.9 62.2 84.5 89.0 51.9 68.2 69.2 71.3
ZeroDL(T-C) 93.5 85.2 50.5 49.3 84.3 66.7 78.3 50.8 74.4 70.3 68.2
IDAS(T-C) 95.3 88.9 45.1 51.6 83.1 56.0 72.2 31.6 43.3 63.0 62.7
Gold(T-C) 94.7 88.5 46.3 49.0 82.9 94.2 98.0 64.5 79.7 77.5 78.6
gemma-2 ZeroDL(C-T) 94.3 88.0 52.6 51.6 74.5 52.8 53.5 50.3 60.0 64.2 60.1
-27b-it IDAS(C-T) 95.3 91.1 42.7 58.3 31.7 56.9 44.0 - - - -
Gold(C-T) 79.8 86.3 48.7 57.3 70.5 89.8 97.1 56.1 65.9 72.4 75.9
ZeroDL(T-C) 94.0 88.3 50.8 50.8 81.8 57.3 57.6 52.7 73.0 67.4 62.7
IDAS(T-C) 95.2 89.2 42.4 55.6 35.0 60.5 49.8 - - - -
Gold(T-C) 95.3 89.6 51.8 55.7 85.4 92.6 98.7 66.9 84.4 80.0 81.0
@@
Table 1: The performance of ZeroDL for text clustering. C-T denotes the prompt order with class information then input text. T-C is the reversed. Bold means the best accuracy in the same prompt order and underline denotes the outperforming cases than baselines (except for Gold label). IDAS with HyperCLOVA X SEED and Gemma-2 is constrained by the maximum 8,192 token length. More baselines are presented in Table 5 of the Appendix.
@@
Neutral Sentiment: This class includes all the sentiment labels that express a neutral sentiment towards the movie or documentary. Examples include (…)
@@
IDAS
Disappointing, Critical, Negative, Positive, Disappointing or Unsatisfying
@@
IDAS
Iran’s Nuclear Program, Oil Prices, Businesses prioritize hardware upgrades in economic recovery, Volcanic activity at Mount St. Helens
@@
Table 2: The example of generated class labels in 5-class sentiment classification (SST-5), and topic classification (AGNews). ZeroDL can generates alternative class labels and its description. Compared with IDAS that has overlap in the generated labels, ZeroDL considers other labels in generation. Furthermore, IDAS’s representative data points fail to generate general topic class labels. Additional examples are in Table 8.
By incorporating the meta-information ($\hat{M}$) into the prompt, we enable the LLM to perform conditioned classification within the clustering context.
4 Experiments
Models
Our primary experiments utilize mistral-7b-instruct-v0.2 Jiang et al. (2023). We further evaluate our approach with several other models, including Qwen-2.5 Team (2024), Gemma-2 Team et al. (2024), Llama-3.1 Meta (2024). Lastly, HyperCLOVA X SEED Yoo et al. (2024) demonstrates the result of a relatively small, instruct-following language model. All the results are averaged over 5 runs.
Baselines
IDAS De Raedt et al. (2023) represents a zero-shot clustering approach. We employ 8 demonstrations as described in the paper, using SBERT Opitz and Frank (2022) as an additional embedder. Final prediction is performed using the same prompt template as our Stage 3. We also include llm2vec BehnamGhader et al. (2024) with K-Means where applicable. This method represents a baseline using (not exactly but) the same backbone computed by embedding-level [^2] Note that IDAS requires an additional embedding model and llm2vec involves additional training steps.. We exclude other few-shot settings from our comparisons as they are not considered fair due to the inherent advantages of leveraging additional information from demonstrations and the subjectivity involved in the few-shot selection.
Setting
Our method aims to effectively cluster data distributions, converging them into a suitable number of classes. We recognize that clustering datasets often exhibit a high number of classes relative to the available data points. To address this, we utilize text classification datasets, which typically offer a larger number of data instances per class than clustering datasets. Moreover, the availability of ground-truth labels facilitates direct comparison between generated and actual labels, enabling qualitative analysis and interpretability.
Datasets
We thus use 6 text classification datasets for clustering: IMDB Maas et al. (2011), SST-2, SST-5 Socher et al. (2013), YelpReivews Zhang et al. (2015) for sentiment classification, and AGNews, DBpedia Lehmann et al. (2015), YahooAnswers Chang et al. (2008) for topic classification. Data information is presented in Table 4 of the Appendix.
Evaluation
Model performance is evaluated based on the highest accuracy achieved across all possible mapping combinations. To determine the predicted class, we leverage the Class n anchor tokens within the LLM outputs. However, LLMs might not directly predict the same classes as the ground-truth labels [^3] For example, Class 0 means Positive in the prediction while Negative is labeled as Class 0 in the ground-truth., so we test all possible mapping combinations and report the best performed one. We split datasets with more than 7 classes (i.e., DBpedia and YahooAnswers) into 2 subsets Front (F) and Back (B) to avoid out-of-memory [^4] The computation is a factorial of the number of classes. after removing the 3 smallest size of classes [^5] CarsAndTransportation, SocialScience, and Sports classes in YahooAnswers dataset..
5 Results
Table 1 presents the performance of our ZeroDL method. Notably, ZeroDL surpasses models provided with ground-truth (Gold) class labels on several datasets. This suggests that ZeroDL may uncover richer or more nuanced class labels within the data compared with pre-defined labels. Our method demonstrates performance comparable to K-Means clustering using LLM embeddings, particularly excelling on datasets with relatively smaller sizes. ZeroDL outperforms IDAS, especially on datasets with more than 2 classes. IDAS often struggles to generate appropriate class labels, particularly when relying on a limited coverage of representative data instances as demonstrations. Furthermore, IDAS is susceptible to the input length limitations of LLMs when the demonstrations are long, as evident in the Gemma results. In contrast, ZeroDL achieves its results through flexible zero-shot prompting without requiring any modifications.
Table 2 shows examples of class labels generated from ZeroDL. These labels often provide richer and more informative explanations [^6] Note that the generation of class description varies depending on datasets and LLMs. compared with the original ground-truth labels. Moreover, ZeroDL can potentially uncover novel classes based on the data, such as Ambiguous Sentiment and Mixed Sentiment [^7] Examples are provided in Table 6 of the Appendix.. This highlights the limitations of pre-defined labels in capturing the full spectrum of sentiment complexity.
Figure 2: Confusion matrix of proposed method ZeroDL in SST-5 (left) and AGNews (right). x-axis and y-axis are generated labels and gold labels, respectively.
Figure 2 illustrates the confusion matrix of the results. The relatively low performance of SST-5, even with gold labels, underscores the inherent difficulty of sentiment classification. In contrast, AGNews demonstrates successful matching with our method. This observation emphasizes the potential value of ZeroDL in scenarios where ground-truth labels are unavailable.
Table 3 investigates the significance of class labels in text clustering tasks. We explore the performance of mistral-7b-instruct-v0.2 using class labels suggested by AutoL Gao et al. (2021) designed for prompt-based model fine-tuning. These labels represent a curated selection. We also investigate the class labels generated by gpt-4.1-mini. The results demonstrate that clustering performance is generally higher when using manually selected labels than using the original dataset labels. This suggests that carefully chosen labels can significantly improve clustering outcomes. In the context, ZeroDL performs particularly well on SST-5, implying that the potential of ZeroDL to capitalize on informative class labels automatically.
ZeroDL involves a trade-off in computational cost. We evaluate the impact of varying input data amounts on performance (see Table 7 in the Appendix). While using only 10% of the data yields reasonable performance, it can lead to inconsistencies in model performance as evidenced by increased standard deviation.